the size of a computer monitor is usually given by the length of its diagonal. a monitors aspect ratio is the ratio of its width to its height. this monitor has a diagonal lenght of 19 inches and an aspect ratio of 5:4 what are the width and height of the monitor? round to the nearest tenth of an inch.
1
Expert's answer
2012-11-12T12:41:18-0500
According to the Pythagoras' theorem we have:

(5x)² + (4x)² = 19² ==>

25x² + 16x² = 361 ==>

41x² = 361 ==>

x = √(361/41) ≈ 3.0 inches.

So, the width of a monitor is about 5x = 5*3.0 = 15.0 inches and its height is about 4*3.0 = 12.0 inches.